OpenStudy (anonymous):

Find parameter a values , in which the equation... will type it... has (5-a+3|a-1|)/4 solutions.

OpenStudy (anonymous):

\[(10a-29)*27^x+(11-3a)*9^x-(7a-17)*3^x+1=0\]

OpenStudy (anonymous):

5

OpenStudy (experimentx):

that equation should have ???

OpenStudy (anonymous):

Find the values of a so that it would have (5-a+3|a-1|)/4 solutions.

jhonyy9 (jhonyy9):

what sign is after 3 ?

OpenStudy (anonymous):

Absolute value

OpenStudy (experimentx):

looks like 5

jhonyy9 (jhonyy9):

this wann to be 3/a ?

OpenStudy (anonymous):

3*|a-1|

OpenStudy (experimentx):

since it's cubic equation, it has 3 roots at most ...

jhonyy9 (jhonyy9):

- so than you will have two cases for Ia-1I 1. when Ia-1I = a-1 2. ---- Ia-1I = - (a-1)

jhonyy9 (jhonyy9):

can you continue it ?

OpenStudy (anonymous):

I don't think so.

OpenStudy (asnaseer):

ErkoT: you need to use the hints that experimentX and jhonyy9 gave you. so we know this is a cubic equation - therefore there are 3 solutions. therefore, we know:\[\frac{5-a+3|a-1|}{4}=3\]solve this for two cases, one where a >= 1 and the other where a < 1.

OpenStudy (anonymous):

Alright. I'll try.